trying to solve this expression with an equal sign
4/(x+3) +2/(x+6) = -6/x^2 +9x + 18
I got that
6/(x+3)(x+6) = -6/(x+3)(x+6)

Answer by stanbon(75887)

trying to solve this expression with an equal sign
4/(x+3) +2/(x+6) = -6/x^2 +9x + 18
I got that
6/(x+3)(x+6) = -6/(x+3)(x+6)
---------------------------------
4/(x+3) +2/(x+6) = -6/x^2 +9x + 18
----
Note: Whenever you have fractions in an equation
you can multiply thru by the lcm to rid of the
denominators.
--------------------------------------------------
Multiply thru by (x^2+9x+18) to get:
4(x+6)+2(x+3) = -6
6x+30 = -6
6x = -36
x = -6
============
Cheers,
Stan H.